On monday Ben drove to work in 35 minutes,On friday he average 9 miles per hour more, and it took him 7 minutes less to get to work. how far in miles does he travel to work.
Answer by stanbon(75887)

On monday Ben drove to work in 35 minutes,On friday he average 9 miles per hour more, and it took him 7 minutes less to get to work. how far in miles does he travel to work.
------------------------
Monday DATA:
time = 35/60 hr ; rate = x mph ; distance = 35x/60 miles
----
Friday DATA:
time = 18/60 hr ; rate = x+9 mph ; distance = rt = (x+9)(18/60) miles
-----------------
Equation:
distance = distance
35x/60 = (x+9)(18/60)
35x = 18x + 9*18
7x = 9*18
x = 23.142857
-----
Distance = 35*23.142857/60 = 13.5 miles
============================================
Cheers,
Stan H.